OpenStudy (anonymous):

Find the x- and y- intercepts of x - 2y = -8. Write them as ordered pairs.

OpenStudy (anonymous):

x-intercept is y = 0. y-intercept is x = 0. So plug these into your equation to get your answer. x - 2y = -8 x - 2(0) = -8 x - 0 = -8 x = -8 So the x-intercept is -8.

OpenStudy (anonymous):

Find the y-intercept: x - 2y = -8 (0) - 2y = -8 -2y = -8 Can you finish this? What do you have to do next?

OpenStudy (anonymous):

You have to divide -8 / -2, what do you get?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Yes, so your y-intercept is 4. You can write this as (-8, 4)

OpenStudy (anonymous):

You can give medals by clicking "Best Response".

OpenStudy (anonymous):

ty so much!

OpenStudy (anonymous):

Np